# Could the neutrino emission of TXS 0506+056 come from the accretion flow of the supermassive black hole?

### v3

@@ 1 Introduction
Active Galactic Nuclei (AGN) are prime candidate sources of the high-energy astrophysical neutrinos. In 2017, IceCube detected a high-energy neutrino event in the direction coincident with the blazar TXS 0506+056, which was found to be flaring at the gamma-ray band (IceCube Collaboration et al., 2018a). A follow-up analysis of archival IceCube neutrino data revealed an earlier outburst of neutrinos from the same source in 2014/2015 without an accompanying flare of gamma rays(IceCube Collaboration et al., 2018b). Later, an independent search for point-like sources in the northern hemisphere using ten years of IceCube data revealed that TXS 0506+056 is coincident with the second hottest-spot of the neutrino event excess (IceCube Collaboration et al., 2022). Since blazars are AGNs with relativistic jets pointing toward our line of sight, the Doppler effect remarkably boosts the flux of blazars received by us. Most of the previous studies ascribe the high-energy neutrino emission of TXS 0506+056 to relativistic protons accelerated in the jet, through either photopion production with the radiation of the jet itself and the external radiation of the surrounding environment (e.g. Murase et al., 2018; Keivani et al., 2018; Gao et al., 2019; Cerruti et al., 2019; Rodrigues et al., 2019; Xue et al., 2019; Zhang et al., 2020; Xue et al., 2021) , or proton-proton collisions with matter of the jet and cloud/star entering the jet (e.g. Sahakyan, 2018; Liu et al., 2019; Banik et al., 2020; Wang et al., 2022). However, the composition of jets of AGNs, whether they are baryon dominated or Poynting flux dominated, is largely unknown. The composition of jets is tightly related to their formation mechanisms. As discussed by Celotti & Blandford (2001), electromagnetically dominated outflows, such as those generated by the extraction of the spin energy of the black hole, will become pair-dominated jets with a low baryonic pollution. In this case, no comic rays and neutrinos would be expected from the AGN jets.
@@
The rest part of the paper is organized as follows. We introduce our model in Section 2 and show the results in Section 3. The discussions are written in Section 4. Finally we summarize the results in Section 5.
@@ 2 Neutrinos from the AGN disk in TXS 0506+056 ?
$\sim 5\times 10^{45}{\rm erg~{}s^{-1}}\left(\frac{\epsilon_{\rm CR}}{0.1}\right)\left(\frac{\dot{m}}{1}\right)\left(\frac{M_{\rm BH}}{3\times 10^{8}M_{\odot}}\right).$
Here $f_{pp,p\gamma}=t_{\rm loss}/t_{pp,p\gamma}$ is the efficiency of $pp$ and $p\gamma$ processes expressed in the ratio between the proton energy loss timescale and the hadronic interaction timescale, which will be calculated in following sections. To explain the observed neutrino luminosity, we require
@@ 2.1 The process of proton acceleration and cooling in the accretion flow
$t_{\rm acc}\approx\frac{\eta r_{\rm L}}{c}$ (5)
where the different mechanisms may be characterized by distinct parameter $\eta$, which could be understood as the particle acceleration efficiency. $r_{\rm L}=E/eB$ is the Larmor radius. The maximum energy of non-thermal proton is determined by the balance among particle acceleration, cooling, and escape processes in the accretion flow. The escape term is common for all components. We consider diffusion and advection (infall to the BH) as the escape processes, whose timescales are estimated to be $t_{\rm diff}\approx R^{2}/D_{\rm R}$ and $t_{\rm fall}\approx R/V_{\rm R}$ respectively, where $D_{\rm R}=D_{\rm R}\approx\eta r_{\rm L}c/3$ is the diffusion coefficient and $V_{\rm R}$ is the radial velocity of the accretion flow. The total escape time is given by $t_{\rm esc}^{-1}=t_{\rm diff}^{-1}+t_{\rm fall}^{-1}$.
@@
The timescale of photomeson process ($p\gamma$) is $t_{p\gamma}\approx 1/(n_{\gamma}\sigma_{p\gamma}\kappa_{p\gamma}c)$, where ${\sigma}_{p\gamma}\approx 5\times 10^{-28}~{}\rm{cm}^{2}$ is the cross section for the photomeson process and $\kappa_{p\gamma}\sim 0.2$ is the inelasticity for $p\gamma$. The Bethe-Heitler energy loss rate is $t_{\rm B-H}\approx 1/(n_{\gamma}\hat{\sigma}_{\rm B-H}c)$, where $\hat{\sigma}_{\rm B-H}\sim 0.8\times 10^{-30}\rm{cm}^{-3}$ is the effective cross section for the Bethe-Heitler process (Murase et al., 2020). The $pp$ cooling timescale is $t_{pp}\approx 1/(n_{p}{\sigma}_{pp}\kappa_{pp}c)$, where ${\sigma}_{pp}\simeq 4\times 10^{-26}\rm{cm}^{2}$ and $\kappa_{\rm pp}\approx 0.5$ are cross section and inelasticity for $pp$ process (Kelner et al., 2006). The proton synchrotron timescale is $t_{p,\rm{syn}}={6\pi m_{p}c}/({\gamma_{p}\sigma_{T}B^{2}})$. The total cooling rate can be given by $t_{\rm cool}^{-1}=t_{p\gamma}^{-1}+t_{pp}^{-1}+t_{\rm syn}^{-1}+t_{\rm B-H}^{-1}$, which is summation of all cooling rate. For high energy protons, the total energy loss rate is $t_{\rm loss}^{-1}=t_{\rm esc}^{-1}+t_{\rm cool}^{-1}$.
@@ 2.2 Proton spectrum
$\frac{d}{dE_{p}}\left(-\frac{E_{p}}{t_{\text{cool}}}N_{p}\right)=\dot{N}_{\rm p,inj}-\frac{N_{p}}{t_{\text{esc}}},$ (7)
@@
$\dot{N}_{p,\text{inj}}=\dot{N}_{0}E_{p}^{-s_{\rm inj}}\exp\left(-\frac{\textit{E}_{{p}}}{\textit{E}_{\textit{p},\text{max}}}\right),$ (8)
@@
Figure 1: The panels, from top to bottom, show the various timescales, the efficiencies of $pp$ and $p\gamma$ interactions, and the neutrino spectrum in the MAD scenario. In panel (c), the observed neutrino spectrum of TXS 0506+056 during the 2014-2015 neutrino flare is also shown(IceCube Collaboration et al., 2018b). We adopt the parameter values of $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{CR}=0.1$, $\epsilon=0.01$, $\eta=300$ and $\dot{m}=10$. In panel (a) and panel (b), the dissipation radius is set as $30R_{g}$, while in panel (c), two radius with $R=30R_{g}$ and $R=60R_{g}$ are assumed.
@@ 3 Neutrino emission from the accretion flow in the MAD and SANE scenarios
A magnetically arrested accretion disc (MAD, see Narayan et al. (2003a); Bisnovatyi-Kogan & Ruzmaikin (1974); Igumenshchev et al. (2003); Tchekhovskoy et al. (2014)) may be present in TXS 0506+056 since magnetohydrodynamical (MHD) simulations have shown that it can launch powerful jets (e.g. Tchekhovskoy et al. (2011)). In this situation, a large-scale poloidal magnetic field prevents gas from accreting continuously at a magnetospheric radius. Around the magnetospheric radius, the gas flow breaks up into a blob-like stream and moves inward by diffusing via magnetic interchanges through the magnetic field. We also consider the standard and normal evolution (SANE) regime of accretion, where the magnetic field that accumulates around the BH is relatively weak.
@@ 3.1 The MAD scenario
In the MAD scenario, the proton number density in the accretion flow is $n_{\rm p,MAD}=\dot{M}/4\pi m_{p}RHV_{\rm R,MAD}$, where $V_{\rm R,MAD}=\epsilon V_{\rm ff}$ is the radial velocity of the accretion flow, $V_{\rm ff}=\sqrt{2GM_{\rm BH}/R}$ is free-fall velocity and $\epsilon\lesssim 0.01$ (Narayan et al., 2003b). Taking $H=R/2$, we have
@@
The magnetic field of MAD can be obtained by equating the magnetic energy density $B_{\rm MAD}^{2}/(8\pi)$ with the gravitational force per unit area of the radially accreting mass $GM_{\rm BH}m_{p}n_{\rm p,MAD}H/R^{2}$ (Hayasaki & Yamazaki, 2019), which is given by
@@
$t_{p\gamma,\rm{MAD}}\approx 1/n_{\gamma}{\sigma}_{p\gamma}K_{p\gamma}c\simeq 10^{4}\,\rm{s}\left(\frac{\textit{n}_{\gamma}}{3\times 10^{13}~{}\rm{cm}^{-3}}\right)^{-1},$ (16)
@@
We show the various timescales in the panel (a) of Fig. 1 for $R=30~{}r_{g}$ and $R=60~{}r_{g}$ with super-Eddington accretion $\dot{m}=10$. We used strict expressions of $p\gamma$ and Bethe-Heitler process
@@
in calculating the relevant timescales, $\gamma_{p}$ is Lorentz factors of protons, $\bar{\epsilon}$ is the photon energy in the proton rest frame, ${dn}/{d\epsilon}$ is the number density of target photons, $\epsilon_{\text{th}}$ is the threshold energy of $p\gamma$ or Bethe-Heitler process, $\sigma$ is cross-section and $\kappa$ is inelasticity for $p\gamma$ or Bethe-Heitler process.
@@
Figure 3: The panels, from top to bottom, show the various timescales, the efficiencies of $pp$ and $p\gamma$ interactions, and the neutrino spectrum in the SANE scenario. We adopt the parameters $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{CR}=0.1$, $\alpha=0.3$, $\beta=10$, $\eta=300$, $\beta=10$ and $\dot{m}=10$. For panel (a) and panel (b), the dissipation radius is set as $30R_{g}$, while in panel (c), two radius with $R=10R_{g}$ and $R=30R_{g}$ are assumed. In panel (c), the observed neutrino spectrum of TXS 0506+056 during the 2014-2015 neutrino flare is also shown.
@@ 3.2 The SANE scenario
$B_{\rm SANE}=\sqrt{\frac{8\pi n_{\rm p,SANE}m_{p}C_{s}^{2}}{\beta}}\sim 2.5\times 10^{3}\rm{G}\left(\frac{\dot{\textit{m}}}{10}\right)^{1/2}$ (21)
@@
From the above timescales, we find that in the SANE scenario, depending on the parameter values, the $pp$, $p\gamma$ and the proton synchrotron emission processes all could become the dominant cooling mechanism for the highest energy protons. If the $pp$ process becomes the dominant mechanism for the proton cooling, by equating the acceleration timescale with the $pp$ cooling timescale, we obtain the maximum proton energy,
@@
Figure 5: The neutrino spectrum with different acceleration efficiency $\eta$ in the MAD scenario. The parameter values used are: $R=60R_{g}$, $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{CR}=0.1$, and $\epsilon=0.01$. The solid line, the dashed line and the dot-dashed line represent $\eta=300$, $\eta=3000$ and $\eta=30000$, respectively. The left panel shows the super-Eddington regime ($\dot{m}=10$) for the 2014-2015 neutrino flare of TCS 0506+056, whereas the right panel shows the sub-Eddington accretion regime ($\dot{m}=0.1$) for the steady-state emission.
Figure 6: Same as Fig.5, but for the SANE scenario. The parameter values used are: $R=30R_{g}$, $M_{\rm BH}=3\times 10^{8}$, $\epsilon_{CR}=0.1$, $\alpha=0.3$, and $\beta=10$.

### v4

@@ 1 Introduction
Active Galactic Nuclei (AGN) are prime candidate sources of the high-energy astrophysical neutrinos. In 2017, IceCube detected a high-energy neutrino event in the direction coincident with the blazar TXS 0506+056, which was found to be flaring at the gamma-ray band (IceCube Collaboration et al., 2018a). A follow-up analysis of archival IceCube neutrino data revealed an earlier outburst of neutrinos from the same source in 2014/2015 without an accompanying flare of gamma rays(IceCube Collaboration et al., 2018b). Later, an independent search for point-like sources in the northern hemisphere using ten years of IceCube data revealed that TXS 0506+056 is coincident with the second hottest-spot of the neutrino event excess (IceCube Collaboration et al., 2022). Since blazars are AGNs with relativistic jets pointing toward our line of sight, the Doppler effect remarkably boosts the flux of blazars received by us. Most of the previous studies ascribe the high-energy neutrino emission of TXS 0506+056 to relativistic protons accelerated in the jet, through either photopion production with the radiation of the jet itself and the external radiation of the surrounding environment (e.g. Murase et al., 2018; Keivani et al., 2018; Gao et al., 2019; Cerruti et al., 2019; Rodrigues et al., 2019; Xue et al., 2019; Zhang et al., 2020; Xue et al., 2021) , or proton-proton collisions with matter of the jet and cloud/star entering the jet (e.g. Sahakyan, 2018; Liu et al., 2019; Banik et al., 2020; Wang et al., 2022). However, the composition of jets of AGNs, whether they are baryon dominated or Poynting flux dominated, is largely unknown. The composition of jets is tightly related to their formation mechanisms. As discussed by Celotti & Blandford (2001), electromagnetically dominated outflows, such as those generated by the extraction of the spin energy of the black hole, are pair-dominated jets with a low baryonic pollution. In this case, no comic rays and neutrinos would be expected from the AGN jets.
@@
The rest part of the paper is organized as follows. We introduce our model in Section 2 and show the results in Section 3. The discussions are given in Section 4. Finally we summarize the results in Section 5.
@@ 2 Neutrinos from the AGN disk in TXS 0506+056 ?
$\sim 5\times 10^{45}{\rm erg~{}s^{-1}}f_{pp,p\gamma}\left(\frac{\epsilon_{\rm CR}}{0.1}\right)\left(\frac{\dot{m}}{1}\right)\left(\frac{M_{\rm BH}}{3\times 10^{8}M_{\odot}}\right).$
Here $f_{pp,p\gamma}=t_{\rm loss}/t_{pp,p\gamma}$ is the efficiency of $pp$ and $p\gamma$ processes expressed in the ratio between the proton energy loss timescale ($t_{\rm loss}$) and the hadronic interaction timescale ($t_{pp,p\gamma}$), which will be calculated in following sections. To explain the observed neutrino luminosity, we require
@@ 2.1 The process of proton acceleration and cooling in the accretion flow
$t_{\rm acc}\approx\frac{\eta r_{\rm L}}{c},$ (5)
where the different particle acceleration mechanisms may be characterized by distinct parameter $\eta$, which could be understood as the particle acceleration efficiency. $r_{\rm L}=E/eB$ is the Larmor radius. The maximum energy of non-thermal proton is determined by the balance among particle acceleration, cooling, and escape processes in the accretion flow. The escape term is common for all components. We consider diffusion and advection (infall to the BH) as the escape processes, whose timescales are estimated to be $t_{\rm diff}\approx R^{2}/D_{\rm R}$ and $t_{\rm fall}\approx R/V_{\rm R}$ respectively, where $D_{\rm R}=D_{\rm R}\approx\eta r_{\rm L}c/3$ is the diffusion coefficient and $V_{\rm R}$ is the radial velocity of the accretion flow. The total escape time is given by $t_{\rm esc}^{-1}=t_{\rm diff}^{-1}+t_{\rm fall}^{-1}$.
@@
The timescale of photomeson process ($p\gamma$) is $t_{p\gamma}\approx 1/(n_{\gamma}\sigma_{p\gamma}\kappa_{p\gamma}c)$, where ${\sigma}_{p\gamma}\approx 5\times 10^{-28}~{}\rm{cm}^{2}$ is the cross section for the photomeson process and $\kappa_{p\gamma}\sim 0.2$ is the inelasticity for $p\gamma$. The Bethe-Heitler energy loss rate is $t_{\rm B-H}\approx 1/(n_{\gamma}\hat{\sigma}_{\rm B-H}c)$, where $\hat{\sigma}_{\rm B-H}\sim 0.8\times 10^{-30}\rm{cm}^{-3}$ is the effective cross section for the Bethe-Heitler process (Murase et al., 2020). The $pp$ cooling timescale is $t_{pp}\approx 1/(n_{p}{\sigma}_{pp}\kappa_{pp}c)$, where ${\sigma}_{pp}\simeq 4\times 10^{-26}\rm{cm}^{2}$ and $\kappa_{pp}\approx 0.5$ are cross section and inelasticity for $pp$ process (Kelner et al., 2006). The proton synchrotron timescale is $t_{p,\rm{syn}}={6\pi m_{p}c}/({\gamma_{p}\sigma_{T}B^{2}})$. The total cooling rate can be given by $t_{\rm cool}^{-1}=t_{p\gamma}^{-1}+t_{pp}^{-1}+t_{\rm syn}^{-1}+t_{\rm B-H}^{-1}$, which is summation of all cooling rate. For high energy protons, the total energy loss rate is $t_{\rm loss}^{-1}=t_{\rm esc}^{-1}+t_{\rm cool}^{-1}$.
@@ 2.2 Proton spectrum
$\frac{d}{dE_{p}}\left(-\frac{E_{p}}{t_{\text{cool}}}N_{p}\right)=\dot{N}_{\textit{p},\text{inj}}-\frac{N_{p}}{t_{\text{esc}}},$ (7)
@@
$\dot{N}_{\textit{p},\text{inj}}=\dot{N}_{0}E_{p}^{-s_{\rm inj}}\exp\left(-\frac{\textit{E}_{{p}}}{\textit{E}_{\textit{p},\text{max}}}\right),$ (8)
@@
Figure 1: The panels, from top to bottom, show the various timescales, the efficiencies of $pp$ and $p\gamma$ interactions, and the neutrino spectrum in the MAD scenario. In panel (c), the observed neutrino spectrum of TXS 0506+056 during the 2014-2015 neutrino flare is also shown (IceCube Collaboration et al., 2018b). We adopt the parameter values of $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{\rm CR}=0.1$, $\epsilon=0.01$, $\eta=300$ and $\dot{m}=10$. In panel (a) and panel (b), the dissipation radius is set as $30R_{g}$, while in panel (c), two radius with $R=30R_{g}$ and $R=60R_{g}$ are assumed.
@@ 3 Neutrino emission from the accretion flow in the MAD and SANE scenarios
A magnetically arrested accretion disc (MAD, see Narayan et al. (2003); Bisnovatyi-Kogan & Ruzmaikin (1974); Igumenshchev et al. (2003); Tchekhovskoy et al. (2014)) may be present in TXS 0506+056 since magnetohydrodynamical (MHD) simulations have shown that it can launch powerful jets (e.g. Tchekhovskoy et al. (2011)). In this situation, a large-scale poloidal magnetic field prevents gas from accreting continuously at a magnetospheric radius. Around the magnetospheric radius, the gas flow breaks up into a blob-like stream and moves inward by diffusing via magnetic interchanges through the magnetic field. We also consider the standard and normal evolution (SANE) regime of accretion, where the magnetic field that accumulates around the BH is relatively weak.
@@ 3.1 The MAD scenario
In the MAD scenario, the proton number density in the accretion flow is $n_{\textit{p},\rm{MAD}}=\dot{M}/4\pi m_{p}RHV_{\rm R,MAD}$, where $V_{\rm R,MAD}=\epsilon V_{\rm ff}$ is the radial velocity of the accretion flow, $V_{\rm ff}=\sqrt{2GM_{\rm BH}/R}$ is free-fall velocity and $\epsilon\lesssim 0.01$ (Narayan et al., 2003). Taking disk height $H=R/2$, we have
@@
The magnetic field of MAD can be obtained by equating the magnetic energy density $B_{\rm MAD}^{2}/(8\pi)$ with the gravitational force per unit area of the radially accreting mass $GM_{\rm BH}m_{p}n_{\textit{p},\text{MAD}}H/R^{2}$ (Hayasaki & Yamazaki, 2019), which is given by
@@
$t_{p\gamma,\rm{MAD}}\approx 1/n_{\gamma}{\sigma}_{p\gamma}\kappa_{p\gamma}c\simeq 10^{4}\,\rm{s}\left(\frac{\textit{n}_{\gamma}}{3\times 10^{13}~{}\rm{cm}^{-3}}\right)^{-1},$ (16)
@@
We show the various timescales in the panel (a) of Fig. 1 for $R=30R_{g}$ and $R=60R_{g}$ with super-Eddington accretion $\dot{m}=10$. We used strict expressions of $p\gamma$ and Bethe-Heitler process
@@
in calculating the relevant timescales, $\gamma_{p}$ is Lorentz factors of protons, $\bar{\epsilon}$ is the photon energy in the proton rest frame, ${dn}/{d\epsilon}$ is the number density of target photons, $\epsilon_{\text{th}}$ is the threshold energy of $p\gamma$ or Bethe-Heitler process, $\sigma$ is cross-section and $\kappa$ is inelasticity for $p\gamma$ or Bethe-Heitler process. Note that the acceleration timescale required for protons to reach the maximum energy of 50 PeV is on the order of $10^{3}-10^{4}{\rm s}$, which is much shorter than the duration of the 2014-2015 flare. Thus, the flare duration does not affect the maximum energy of accelerated protons.
@@
Figure 3: The panels, from top to bottom, show the various timescales, the efficiencies of $pp$ and $p\gamma$ interactions, and the neutrino spectrum in the SANE scenario. We adopt the parameters $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{\rm CR}=0.1$, $\alpha=0.3$, $\beta=10$, $\eta=300$, $\beta=10$ and $\dot{m}=10$. For panel (a) and panel (b), the dissipation radius is set as $30R_{g}$, while in panel (c), two radius with $R=10R_{g}$ and $R=30R_{g}$ are assumed. In panel (c), the observed neutrino spectrum of TXS 0506+056 during the 2014-2015 neutrino flare is also shown.
@@ 3.2 The SANE scenario
$B_{\rm SANE}=\sqrt{\frac{8\pi n_{p,{\rm SANE}}m_{p}C_{s}^{2}}{\beta}}\sim 2.5\times 10^{3}\rm{G}\left(\frac{\dot{\textit{m}}}{10}\right)^{1/2}$ (21)
@@
From the timescales above, we find that in the SANE scenario, depending on the parameter values, the $pp$, $p\gamma$ and the proton synchrotron emission processes all could become the dominant cooling mechanism for the highest energy protons. If the $pp$ process becomes the dominant mechanism for the proton cooling, by equating the acceleration timescale with the $pp$ cooling timescale, we obtain the maximum proton energy,
@@
Figure 5: The neutrino spectrum with different acceleration efficiency $\eta$ in the MAD scenario. The parameter values used are: $R=60R_{g}$, $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{\rm CR}=0.1$, and $\epsilon=0.01$. The solid line, the dashed line and the dot-dashed line represent $\eta=300$, $\eta=3000$ and $\eta=30000$, respectively. The left panel shows the super-Eddington regime ($\dot{m}=10$) for the 2014-2015 neutrino flare of TXS 0506+056, whereas the right panel shows the sub-Eddington accretion regime ($\dot{m}=0.1$) for the steady-state emission.
Figure 6: Same as Fig. 5, but for the SANE scenario. The parameter values used are: $R=30R_{g}$, $M_{\rm BH}=3\times 10^{8}M_{\odot}$, $\epsilon_{\rm CR}=0.1$, $\alpha=0.3$, and $\beta=10$.
@@ 4 Discussions
GeV-TeV gamma-rays are observed from TXS 0506+056. Based on the $\gamma\gamma$ optical depth calculations by (Murase et al., 2020), $\gamma\gamma$ absorption in the compact corona and accretion disk can effectively suppress GeV-TeV gamma-ray emission from hadronic processes (i.e. the so-called hidden gamma-ray source). The observed GeV-TeV gamma rays from TXS 0506+056 may then arise from the jets, as usually assumed for GeV-TeV blazars (Madejski & Sikora, 2016; Gao et al., 2019; Liu et al., 2023). These emissions are produced through synchrotron self-Compton (SSC) and/or external Compton (EC) processes of relativistic leptons. Being far from the compact corona/disk, these high-energy gamma-rays can avoid absorption and become detectable.
Most previous studies have suggested that the high-energy neutrinos from TXS 0506+056 originate from the relativistic jet in a region with a radius much larger than that of the accretion flow (Gao et al., 2019; Keivani et al., 2018; Cerruti et al., 2019). One way to distinguish between our accretion flow model and the jet model is through the cascade emission induced by the absorbed high-energy gamma-rays accompanying the neutrino production. The compact size of the neutrino production region in the accretion flow model leads to strong absorption of the high-energy gamma-rays and the absorbed energy is transferred to MeV emission through the electromagnetic cascades(Murase et al., 2020), Future mission such as AMEGO-X (Caputo et al., 2022) and eASTROGAM (De Angelis et al., 2017) can detect the MeV emission and test the accretion flow model. In addition, our model predicts that mis-aligned blazars (i.e., radio galaxies), where the jets do not point toward us, can also produce high-energy neutrinos through the accretion flow. Future observations by next-generation neutrino telescopes, such as IceCube-Gen2 (IceCube-Gen2 Collaboration, 2023) and HUNT (Huang et al., 2024) can test this.